OpenStudy (anonymous):

how do we solve theta in this? sin(theta)cos(theta)=1/4

OpenStudy (loser66):

do you know that 2sin \(\theta\) cos\(\theta\)= sin 2\(\theta\)?

OpenStudy (anonymous):

nope.

OpenStudy (loser66):

so, now you know it, then solve from then. Don't ask me "how" because I don't know.

OpenStudy (anonymous):

how? haha. i don't know either.

OpenStudy (loser66):

I told you, don't ask me "how" hahaha.... I can predict that question. XD

OpenStudy (akashdeepdeb):

Do you know sin(x+y) ?

OpenStudy (akashdeepdeb):

@student12345 ? sin2theta = 2.sintheta.cos theta!

OpenStudy (anonymous):

how? @AkashdeepDeb can you do the procedure step by step? i feel confused.

OpenStudy (akashdeepdeb):

Okay so there is one thing you'd HAVE to learn and that is sin(x+y) = sinx.cosy + siny.cosx Alright?

OpenStudy (anonymous):

yes.

OpenStudy (akashdeepdeb):

Now what is \[\sin 2\theta = \sin (\theta + \theta) = \sin \theta.\cos \theta + \sin \theta.\cos \theta = 2\sin \theta.\cos \theta\] According to the law above! Getting this? :)

OpenStudy (anonymous):

no. :s

OpenStudy (akashdeepdeb):

sin(x+y) = sinx.cosy + siny.cosx sin(θ+θ)=sinθ.cosθ+sinθ.cosθ=2sinθ.cosθ x=y=\[\theta\] Understood? :)

OpenStudy (anonymous):

yes yes. but how could we rely it to sin(theta)cos(theta)=1/4 ?

OpenStudy (akashdeepdeb):

Okay so now Let us take it like this! 4.sinθ.cosθ = 1 Right? :)

OpenStudy (akashdeepdeb):

or 2(2.sinθ.cosθ) = 1 or sin2θ = 1/2 [Inverse of that previous statement] So sin2θ = sin 30 = 1/2 Or 2θ = 30 or θ = 15' Understood now? :D

OpenStudy (anonymous):

we have formulae: \[\sin(x+y) = sinx.cosy + cosx.siny\] \[\sin(x-y) = sinx.cosy - cosx.siny\] \[\cos(x+y) = cosx.cosy - sinx.siny\] \[\cos(x-y) = cosx.cosy + sinx.siny\] \[\tan(x+y) = \frac{ tanx + tany }{ 1 - tanx.tany }\] \[\tan(x-y) = \frac{ tanx - tany }{ 1 + tanx.tany }\] remember these

OpenStudy (anonymous):

@student12345 can u now give me 2θ's of the three trig functions, cos , sin and tan from above equations?

OpenStudy (anonymous):

the 2θ's are important for u , u need to by-heart them after you learned how they actually derived.